if a child pours 210ml juice in a glass and then drinks one seventh of it, after which he adds another 30% of what is left in the glass to it, how much juice is in the glass


1
Expert's answer
2022-02-09T06:29:04-0500

Initially: 210 mL. Drinks 1/7 of it, what left is


"V_1=210·(1-1/7)=180\text{ mL}."


Then the child adds 30% of what is left in the glass to the 180 mL:


"V_2=180(1+30/100)=234\text{ mL}."